OpenStudy (anonymous):

how do I integrate: ln((x+2)/(x+1)) ?

OpenStudy (phi):

use ln( a/b) = ln(a) - ln(b) to break up the problem

OpenStudy (anonymous):

I went from that to the equation I have written above.

OpenStudy (amistre64):

why would you combine it to make it harder? phi is right, split it up

OpenStudy (phi):

integration by parts. \[\int\limits u dv= u v- \int\limits v du\] for the 1st integral ln(x+2) dx u= ln(x+2) dv = dx

OpenStudy (phi):

now find du = d/dx ln(x+2) = 1/(x+2) dx and v = integral dx = x

OpenStudy (phi):

can you finish?

OpenStudy (anonymous):

No I cant

OpenStudy (anonymous):

http://openstudy.com/users/borgerkrig#/updates/4f565507e4b0862cfd074d0b Can you please check this link. Maybe my question seems out of the blue, because it's connected to the question in the link.